A cashier has 2 times as many nickels as quarters, 15 more dimes than quarters that worth $3.75 in total. Find how many dimes she has.
Answer by nerdybill(7384)

A cashier has 2 times as many nickels as quarters, 15 more dimes than quarters that worth $3.75 in total. Find how many dimes she has.
.
Let q = number of quarters
then from: "2 times as many nickels as quarters"
2q = number of nickels
and from: "15 more dimes than quarters"
q+15 = number of dimes
.
Because the cashier has a total of $3.75:
3.75 = .25q + .10(q+15) + .05(2q)
3.75 = .25q + .10q + 1.5 + .1q
3.75 = .45q + 1.5
2.25 = .45q
5 = q (number of quarters)
.
Number of dimes:
q+15 = 5+15 = 20 (number of dimes)